OpenStudy (anonymous):

Solve for n. 22 + n - 7 = 8n - 3n + n n = 5 4/5 n = 5 n = 3 n = 2

OpenStudy (anonymous):

15+n=6n 5n=15 n=3

Parth (parthkohli):

-> 22+n-7 = 6n => 22-7 = 5n => 15 = 5n => n = 3